OpenStudy (anonymous):

plz help evaluate (3/2)^-3

OpenStudy (ibbutibbu.):

are you doing some kind of exam?

OpenStudy (anonymous):

na summer homework trying to catch up

OpenStudy (ibbutibbu.):

ok then

OpenStudy (ibbutibbu.):

3/2 is 1.5 so 1.5^-3... im pretty sure that you use\[\frac{1 }{x^y }\]

OpenStudy (the_fizicx99):

Essentially, you would flip it around. You already have a fraction so this becomes \(\ \sf (\dfrac{3}{2})^{-3} \implies (\dfrac{2}{3})^3 = \dfrac{2^3}{3^3} = \dfrac{8}{27}\)

OpenStudy (anonymous):

ooooooo ok thank you

OpenStudy (the_fizicx99):

Yw ^_~